OpenStudy (anonymous):

If f(x) = x2 – 1 and g(x) = 2x – 3, what is the domain of (f*g)(x)

OpenStudy (solomonzelman):

you are multiplying a polynomial times another polynomial. This will give you a polynomial as well.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

You know what polynomial is, correct?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Ok, will you agree with me (if you don't understand why it is so, then ask! but will you agree) that a polynomial is DEFINED for ALL values of x?

OpenStudy (anonymous):

why is it so? haha

OpenStudy (solomonzelman):

Ok, i will post a percise and short definition of a polynomial. "Polynomial" is a function that can have the terms of: 1) Constants (some real numbers) 2) x raised to a power of 1, 2, 3 and on (but the power has to be a whole number) (and the terms can have any real coefficients)

OpenStudy (solomonzelman):

i will do some examples

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

polynomials. examples. y=0 y=-45 y=√3 (or any function y=c where c is a constant) y=-4x y=3x+2 y=x-9 (or any function in a form of y=mx or y=mx+b) y=3x² y=-5x²+9 y=x²+8 y=x²-7x-7 and so on....

OpenStudy (anonymous):

ok im pretty sure i get it

OpenStudy (solomonzelman):

it (a polynomial) terms can be made up and can consist of any terms as long as the terms have x^(whole number) and other real numbers ok got the definition

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

So, really try to think of any polynomial that can have an UNDEFINED X-VALUE. what I mean by "UNDEFINED X-VALUE" is that if you plug this value for x into the function, then you will get an undefined/indeterminate output for y in polynomials this "UNDEFINED X-VALUE" never exists

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

Because if your x's have only whole number powers such as: x\(^1\), x\(^2\), x\(^3\), x\(^4\) then you will get a valid result NO MATTER WHAT you plug in for x.

OpenStudy (anonymous):

okay makes sense

OpenStudy (solomonzelman):

now agin, when you multiply two functions where both of these functions are polynomials (i.e. whole number powers of x like 1, 2, 3 and on with or without constants) then you get after multiplying these two (or any number of) polynomial functions - you will get - a polynomial function (whole number powers of x like 1, 2, 3 and on with or without constants)

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

polynomial • polynomial = polynomial it is good to remember, even if you don't fully comprehend the deepness behind it (you will get it at some time)

OpenStudy (solomonzelman):

now, if your function f(x)=x²-1 and g(x)=2x-3 polynomial?

OpenStudy (solomonzelman):

are the polynomials?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

they are

OpenStudy (solomonzelman):

so the product of them (when you multiply f • g) will also give you a what?

OpenStudy (anonymous):

polynomial

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

And again, ANY POLYNOMIAL has a domain of ALL VALUES OF X. (in interval notation: Domain: (-∞, +∞) )

OpenStudy (solomonzelman):

So will what be the domain of your new function be? (we know your new function is going to be a polynomial -as we said b4)

OpenStudy (solomonzelman):

if you got questions, you can always ask ....

OpenStudy (anonymous):

im following as of right now :)

OpenStudy (solomonzelman):

we know your function is a polynomial (correct?) So can you give me the domain of your new function, or not yet?

OpenStudy (anonymous):

(-∞, +∞)?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

the ∞ (or +∞) symbol means infinity and -∞ means negative infinity --> (same stength but going the negative way)

OpenStudy (anonymous):

okay i remember that one

OpenStudy (solomonzelman):

(-∞, +∞) means from negative infinity to positive infinity and that is: every single number that there can possible be

OpenStudy (anonymous):

i wish you were my math teacher haha

OpenStudy (solomonzelman):

that is the domain of your new polynomail function, or of any polynomial function.....

OpenStudy (solomonzelman):

i am just some 19 yrs old dude....

OpenStudy (anonymous):

YOUR ONLY 19!!!!

OpenStudy (solomonzelman):

Math is just something I like:) In any case, do you have questions regarding this question?

OpenStudy (anonymous):

nope i got it :) ty so much

OpenStudy (solomonzelman):

k, yw